OpenStudy (anonymous):

Please Help! Yuri solved the equation below incorrectly. Explain the step where he went wrong. 15 x b + 1= 30 Step 1: 15 x b/15 + 1 = 30/15 Step 2: b+1=2 Step 3: b + 1-1=2-1 Step 4:b=1

OpenStudy (anonymous):

Wait wht? Lol 15b+1=30 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Step 1 is wrong because you gotta divide the whole expression by 15 and not 15b. You have to divide 1 by 15

OpenStudy (anonymous):

Psh this is easy but if that's correct then this is what you do. 15b + 1 = 30 -1 -1 15b = 29 divide by 15 on both sides. b = 1.93333 If you can convert that to a fraction, then go for it.

OpenStudy (anonymous):

Thanks Guys